Selected T-accounts of Company are given below for the just completed year:

Raw Materials
Bal. 1/1 30,000 Credits ?
Debits 150,000
Bal. 12/31 40,000
Manufacturing Overhead
Debits 189,500 Credits ?
Work in Process
Bal. 1/1 35,000 Credits 515,000
Direct materials 105,000
Direct labor 195,000
Overhead 214,500
Bal. 12/31 ?
Factory Wages Payable
Debits 215,000 Bal. 1/1 16,500
Credits 210,000
Bal. 12/31 11,500
Finished Goods
Bal. 1/1 55,000 Credits ?
Debits ?
Bal. 12/31 82,500
Cost of Goods Sold
Debits ?

Required:

1. What was the cost of raw materials used in production during the year?

2. How much of the materials in (1) above consisted of indirect materials?

3. How much of the factory labor cost for the year consisted of indirect labor?

4. What was the cost of goods manufactured for the year?

5. What was the unadjusted cost of goods sold for the year? Do not include any underapplied or overapplied overhead in your answer.

6. If overhead is applied to production on the basis of direct labor cost, what predetermined overhead rate was in effect during the year?

7. Was manufacturing overhead underapplied or overapplied? By how much?

8. Compute the ending balance in Work in Process. Assume that this balance consists entirely of goods started during the year. If $11,750 of this balance is direct labor cost, how much of it is direct materials cost? Applied overhead cost?
Answer #1

1) Raw material used = 30000+150000-40000 = 140000

2) Indirect material = 140000-105000 = 35000

3) Indirect material = 210000-195000 = 15000

4) Cost of goods manufactured = 515000

5) Unadjusted cost of goods sold = 55000+515000-82500 = 487500

6) Overhead rate = 214500/195000 = 110% of labor cost

7) Applied overhead = 214500

Actual overhead = 189500

Over applied overhead = 214500-189500 = 25000

8) Ending WIP = 34500

Applied overhead = 11750*1.1 = 12925

Direct material = 34500-11750-12925 = 9825